2x+4y+3z=10
3x - y + 6z = 15
5x + 2y - z = 25
three equation three unknown
what is x, y, and z?

Answer by Boreal(15235)

2x+4y+3z=10
3x - y + 6z = 15
5x + 2y - z = 25
Cramer's Rule
Make Matrix by 2nd Matrix, scroll to edit and hit 3 enter 3 enter.
Input the data
go to quit.
Go to matrix, then math, and 1 det. Go back to matrix and select the matrix and enter it.
Hit enter.
The determinant for the denominator is
2 4 3
3-1 6
5 2 -1
That equals 143
Dx
10 4 3
15 -1 6
25 2 -1
That equals 715, and 715/143=5 for x.
Dy
2 10 3
3 15 6
5 25 -1
That is 0.
Dz
2 4 10
3 -1 15
5 2 25
That is also 0.
(5,0,0)
Works in the first, second and third.